Find the sum of squares of all numbers x such that both expressions (x^2) + 6x and x + (1/x) are
integer numbers.
1
Expert's answer
2020-11-29T19:07:29-0500

Let "x^2+6x=m" and "x+\dfrac{1}{x}=n," where "m" and "n" are integer numbers.

If "x" ≠"\ 0" , then

"\begin{cases}
 x^2+6x=m\\
 x+\dfrac{1}{x}=n
\end{cases}<=>\begin{cases}
 x^2+6x-m=0\\
 x^2-nx+1=0
\end{cases}"


"<=>\begin{cases}
 nx-1+6x-m=0\\
 x^2-nx+1=0
\end{cases}"


"(n+6)x=m+1"

If "n=-6" , then "m=-1" and system of equations "(1)" is equivalent to "x^2+6x+1=0" .


"x_{1,2}=\dfrac{-6\pm \sqrt{36-4}}{2}=-3\pm \sqrt{8},"

and "(x_1)^2+(x_2)^2=9+12\sqrt{2}+8+9-12\sqrt{2}+8=34"


If "n" ≠ "-6" , then "x=\dfrac{m+1}{n+6}" is a rational number.

"x" is one of the roots of "x^2-nx+1=0" . The roots are "x_{1,2}=\dfrac{n\pm\sqrt{n^2-4}}{2}." 

We know, that "x" is rational, therefore "\sqrt{n^2-4}" is integer, because square root of integer number is either integer or irrational. If it was irrational, then fraction "\dfrac{n\pm\sqrt{n^2-4}}{2}" wouldn’t be a rational number.

So, "n^2-4=k^2" , where "k" is integer.

"n^2-k^2=4\quad \Leftrightarrow \quad (n-k)(n+k)=4"

"n-k" and "n+k" are both even numbers

(if they are odd, then their product is odd; if one number is odd and another one is even, then their sum is odd, but it is equal to "2n" )

We have two cases:

1) "\begin{cases}
n-k=2
\\
n+k=2
\end{cases}" and then "n=2""k=0" .

2) "\begin{cases}
n-k=-2
\\
n+k=-2
\end{cases}" and then "n=-2""k=0."

If "n=2" : "x^2-2x+1=(x-1)^2=0" and "x=1" . If this case "x^2+6x" and "x+\dfrac{1}{x}" are integers.

If "n=-2" : "x^2+2x+1=(x+1)^2=0" and "x=-1" . If this case "x^2+6x" and "x+\dfrac{1}{x}" are integers.


So, the sum of squares is "34+1+1=36"